a bullet moving with a velocityof 100m/s can just penetrate two planks of equal thickness. thenumber of such planks penetrated by the same bullet when the velocity is doubled will be 
  1. 4
    1. 6
    2. 8
    3. 10

Grade:11

3 Answers

Shubham Singh
8 years ago
8
velocity is doubled means kinetic energy is made 4 times
Bipin Mehta
6 years ago
3rd equation of motion can be be applied: v^2 = u^2 + 2as, where acceleration is -ve. Hence v^2 = u^2 – 2as.
v, final velocity = 0. Therefore, u^2 = 2as.
i.e. s is proportional to u^2. → s1÷ s2 = u1^2 ÷ u2^2
Therefore, s2 = (u2^2 ÷ u1^2) x s1 = (200÷100)^2 x 2 (planks) = 8 planks
Answer is 8
Saroj Paudel
5 years ago
 Let the velocity be V and thickness be h ( initial) H (final) 
Now
V^2=2gh.   h= thickness
Againa 
(2v)^2=2gH
4v^2=2gH
4*2gh=2gH
4h=H
H=8 .  h=2plancks